Imagine that we took a large sample of respondents and measured their education. The sample mean =16 years and the sample standard variation = 1.5 : using that information, what percentage of people have more than 20 years of education. Show work.
Answer #1

mean = 16 , s = 1.5
Using central limit theorem,

z =(x - mean)/SE

P(x> 20)

=P(z> (20 -16)/1.5)
= P(z> 2.67)

= 1 - P(z< 2.67)
= 1 - 0.9962
= 0.0038